For i =square root −1, if 3i(2 + 5i) = x + 6i,then x = ?
A. –15
B. 5
C. 5i
D. 15i
E. 27i

Answer by jim_thompson5910(35256)

3i(2 + 5i) = x + 6i


6i + 15i^2 = x + 6i


6i + 15(-1) = x + 6i


6i - 15 = x + 6i


- 15 = x


x = -15


So the answer is choice A